

********************************************************************
UARS WEEKLY STATUS REPORT - 04 FEB 2000
********************************************************************
***********
SOLSTICE
***********
Daylight only operations began for SOLSTICE on Thursday 2/3 and will
continue through Monday 2/7. During this time we will only gather solar
data. Otherwise, SOLSTICE has been operating normally, collecting solar
data on all available orbital tracking periods and stellar calibration
observations whenever possible. The reduced tape recorder operations mode
has not prevented us from gathering enough data to meet all the SOLSTICE
science objectives.
***********
SUSIM
***********
Nothing received.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM/AXIS was activate from January 15 through January 24 in the half
AXIS mode. Only AXIS-1 is active, producing energy deposition from electrons
inferred from the measured X-rays from eight pixels on one side of the
spacecraft.
The PEM spectrometers continue to be active under PMON control*.
While under PMON control, the PEM spectrometers computed energy deposition
from electrons and protons measured in-situ by the particle spectrometers.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
MLS was taken out of 'standby' mode on 31 January 2000 and operated for the
first time since 27 July 1999. The objective of this operation is to obtain
information on Arctic vortex ozone chemistry. MLS is being operated in a mode
where limb scans are attempted at ~50N-80N latitudes on the "day" side of the
orbit.
The MLS electronics now on (those needed to support the 205 GHz radiometer and
measurements of ClO, HNO3, O3, UTH, CH3CN) have operated perfectly. No slips
of the antenna scan mechanism (which had plagued MLS operations before putting
the instrument in 'standby') have occurred so far in this session.
The incoming data are now being processed, but a problem has occurred with the
Level 1 processing at the CDHF which we are currently attempting to resolve.
MLS is scheduled to continue observations through 6 February.
The MLS "DAAC Quality Document" for version 5 data has been completed and is
being (or has been) put on the DAAC for distribution with the Version 5 data.
This document is also available on the MLS web site (http://mls.jpl.nasa.gov)
by clicking on "UARS MLS Data" and following pointers there.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Data re-processing and analysis are continuing with WINDII, which
is next scheduled to operate March 5 - 16, 2000.
***********
MPG
***********
The MPG participated in the February UARS Mission Status Review
meeting on 1 February 2000. At this meeting it was decided to
continue the FOT's manually scheduling of NBTR B so that the
maximum number of HALOE observations can be received. HALOE
observations (sunrise and sunset) which will occur between real-
time TDRSS contacts are being recorded on NBTR B in its pass 2
mode. Development of a scenario for 3-TDRSS operation is on-
going.
The schedule for instrument operations for the yaw period 15
February 2000 to 21 March 2000 was distributed and also posted
on the UARS Science homepage http://umpgal.gsfc.nasa.gov/.
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF continues to closely monitor production jobs CPU and Wallclock
usage and notify PI's of any discrepancies encountered in the processing as
a result of the UARS Spacecraft Anomaly.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
1 day HRDI (4.4) PROD (L1-L3) (TEST) (1R-1C=0L)
29 days WIN (5.11A) PROD (L1-L3) (520R-373C-14F=133L)
1 day SOL (4.62) PROD (L1-L3) (TEST) (1R-1C=0L)
CDHF REHOST:
- Testing BFX interface (Level0,NRT).
- Prepare to test New SUSIM patch delivery.
- Continue to test SOLSTICE V4.62 Engineering Software on the
Rehost System.
- Preparing to test HRDI V4.4 on the Rehost System with a different day.
- Plans to test WINDII V5.11A on the Rehost System.
- Data Migration continues.
- Contacted the DAAC regarding Transl8 status for the new alpha
systems, so that the DAAC interface can be tested.
CDHF will load December 1999 current day DAAC data in February 2000.
***********
FOT
***********
REPORTING PERIOD
This report covers 22 January, 2000 (Orbit 45709, GMT Day 022) through 28
January 2000 (Orbit 45813, GMT Day 028).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HRDI 023/1702 45734 Stowed
HRDI 024/0726 45743 Stowed
PEM 024/2248 45753 AXIS-1 Off
HRDI 025/1416 45763 Stowed
HRDI 027/1056 45790 Stowed
The beta angle ranged between the angles of 30.7 degrees to 59.4 degrees for
this report period. The beta angle is now increasing toward a maximum angle of
76.0 degrees on 27 January 1999 (DOY 027).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.0 to 1.3 degrees C. Additional battery performance data for each
battery on 21 January 2000 and 28 January 2000 are provided in Appendix D.
The clock error ranged from 10.3 to -10.1 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% OUT-OF-LIMITS = 04
% ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% HRDI = 04
% FOT = 03
% PEM = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 01
Three (3) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
- Transponder A: 99/358 - 2287.496855
- Transponder B: 99/365 - 2287.499111
GROUND SYSTEM OPERATION
One (1) operational support problems occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/027 show the observatory data loss to be 565 hours, 09
minutes, 54.761 seconds (an increase of 53 minutes 43.936 seconds since the
last report period).
The increase of 09 minutes 54.761 seconds occurred over two (2) days:
- 52 minutes 56.448 seconds on Day 021 due to an unspecified problem.
- 00 minutes 47.488 seconds on Day 026 due to an unspecified problem.
These calculations are under repair. They are being recalculated from the date
where PASS 2 operations were scheduled exclusively.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
No significant satellite operations to report.
MEETINGS
No meetings were held during this report period.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for February 15, 2000 (DOY 046). This
maneuver will be a Reverse to Forward Flight maneuver (Southern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 28 January 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
45734 023/1702 00-026 HRDI SAFEHOLD; Barrel Temp >45 deg.
45743 024/0544 00-028 OBC Clk Adj 20.6 to 41.2 msec/day
45743 024/0726 00-026 HRDI SAFEHOLD: Barrel Temp >45 deg.
45753 024/2248 00-029 PEM AXIS1 Power OFF
45763 025/1416 00-026 HRDI HRDI Safehold - Hi Barrel temp
45787 027/0513 00-027 OBC Clk Adj fr 41.2 to 20.6 msec/day
45790 027/1056 00-026 HRDI HRDI Safehold - Hi Barrel temp
45806 028/1059 00-028 OBC Clk Adj fr 20.6 to 41.2 msec/day
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 28 January 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ -----
PEAXTHERM6E Y-HI PEM AXIS THERMISTOR 6 continuous
HRPTSBARTMP Y-LO HRDI TELE. BARREL TEMP 45724-31, 37, 39-45,
58, 65, 45803-05
HRPTSBARTMP R-LO HRDI TELE. BARREL TEMP 45732-36, 45763, 45806
QTZEPSIF1/2 Y-LO PEM ZEPS IF TEMPS 45733-40
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
No Alerts to report this report period.
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 28 January 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
45743 024/05:47 (24 Jan 00) from 20.6 to 41.2 msec/day
45787 027/05:15 (27 Jan 00) from 41.2 to 20.6 msec/day
45806 028/11:03 (28 Jan 00) from 20.6 to 41.2 msec/day
_______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 28 January 2000
21 Jan 2000 - GMT Day 021 Beta = 59.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ----- ----- ------
1 -61.6/-89.6 -3.05/-4.94 +0.0/+0.0 16.2 0.000 0.0
2 +11.2/-22.4 +4.51/+2.60 +24.0/-15.2 26.4 1.040 17.6
3 +33.6/-11.2 +3.23/+1.66 +23.2/-14.4 26.4 1.038 17.3
28 Jan 2000 - GMT Day 028 Beta = 74.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ----- ------- ------
1 -67.2/-89.6 -2.42/-3.68 +0.0/+0.0 16.2 0.000 0.0
2 +0.0/-33.6 +2.92/+1.66 +16.8/-14.8 27.5 1.057 14.0
3 +44.8/-5.6 +1.97/+1.03 +16.8/-14.4 27.5 1.061 14.5
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ----- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 28 January 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-001 FOT 45803 028/0834 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
No new reports generated
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
No new reports generated
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-064 FOT 44135 10/8-9/99 N/A HRDI Cal Lamp Errors
99-071 FOT 44398 299/1212 21512 TPF Not Receiving 512 PB
00-001 FOT See list January 14170 Generic Late Acquisition
SEVNTFW00/04